OpenStudy (anonymous):

how to find this limit: lim (xˆ2-1)/(x-1) when x goes to 1?

OpenStudy (anonymous):

\[\lim_{x \rightarrow 1}(x ^{2}-1)/x-1\]

OpenStudy (anonymous):

\[\lim_{x \rightarrow 1} (x ^{2} -1)/(x-1) = \lim_{x \rightarrow 1} [(x -1)(x+1)]/(x-1) = \lim_{x \rightarrow 1}(x+1) = 2\]

OpenStudy (anonymous):

remember there is a hole at F(1)

OpenStudy (anonymous):

remember x never equal to 1 but all ways in deleted neighbourhood 1

OpenStudy (anonymous):

2 anavega yeah, we have a hole, but we also have a problem 0/0 =) but we can do the dividing 'cause we have lim when we work with lim we have extra privileges

OpenStudy (anonymous):

remember the concept of lim works only on deleted neighbour of a \[\left( a-\xi,a+\xi \right)\ \left\{ a \right\}\]

OpenStudy (anonymous):

a is not there in the set

OpenStudy (anonymous):

so cannot get 0/o form

OpenStudy (anonymous):

okay, you know about the Hopital's rule? http://en.wikipedia.org/wiki/L%27H%C3%B4pital%27s_rule it say \[\lim_{x \rightarrow c} f(x)/g(x) = [0/0] = \lim_{x \rightarrow c} f(x) \prime/g(x) \prime\] if you do that we should take \[\lim_{x \rightarrow 1} (x ^{2}-1)/(x-1) =\lim_{x \rightarrow 1} 2x = 2\] isn't it?

OpenStudy (anonymous):

it is not =0/=0 cheak it is\[\rightarrow0/\rightarrow0\] =0/=0 is undefined but \[\rightarrow0/\rightarrow0\] is indeterminate form

OpenStudy (anonymous):

In calculus, Hopital's rule p (also called Bernoulli's rule) uses derivatives to help evaluate limits involving indeterminate forms.

OpenStudy (anonymous):

ha then it is \[\epsilon/\epsilon\] epsilon\[\epsilon \rightarrow0\]

OpenStudy (anonymous):

so we cancel the \[\epsilon\]

OpenStudy (anonymous):

you mean that \[\lim_{\epsilon \rightarrow 0} \epsilon/\epsilon = 1\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yeah, you are right, 'cause in this case we have indeterminate form like you say previously \[\lim_{x \rightarrow 0} x/x = [0/0] use Hopital's rule = \lim_{x \rightarrow 0} 1/1 = 1\]

OpenStudy (anonymous):

you can use the l'Hôpital's rule。 lim (xˆ2-1)/(x-1) it is the form of 0/0, because if we make x=1, so x^2-1=0, x-1=0, so it is the form of 0/0 so, at that time we can use l'Hôpital's rule to differentiate both denominator and numerator. so lim (xˆ2-1)/(x-1) =d(x^2-1)/d(x-1)=2x=2

OpenStudy (anonymous):

L'hopital's Rule is okay, but here there is an even easier answer x^2-1=(x+1)(x-1) by factoring so: (x^2-1)/(x-1)=x+1 Then, when x goes to 1 it is clear that f(x) goes to 2.